The goose has a mass of 15.9 lb (pounds) and is flying at 8.10 miles/h (miles per hour). What is the kinetic energy of the goose in joules? Enter your answer numerically in joules.
Answer #1

As 1 pound (lb) = 0.454 kg

So, 15.9 lb = 15.9 x 0.454 = 7.2186 kg

also velocity 1 mile = 1609.344 m

so , 8.10 miles = 8.10 x 1609.344 = 13035.6864 m

so, 8.10 miles /h = 13035.6864 / ( 60 x 60 )

= 3.621 m /s

so that as kinetic energy is given as ,

K = mv^2 / 2 = [ 7.2186 x ( 3.621 ) ^2 ] / 2

K = 94.6476 / 2 = 47.3238 J